A truck of width 6 m is moving on a straight road with a speed of 5m/s.Aperson at a distance of 8m from the truck wants to cross the road safely with minimum possible speed.Find the magnitude and direction of velocity ?Please give the full solution


 


 


Grade:12th Pass

1 Answers

Praneetha reddy
11 years ago

here for truck

speed=5m/s,distance to be covered = 8m

therefore time taken t=distance / speed

=8/5=1.6 sec

now for man

distance to be covered s=6m,time=1.6sec

therefore

speed = distance/time

=6/1.6

=3.75 m/s is the req speed perpendiculur to the direction of truck